Rene, a chemist, needs to produce a 400 ml of 60 percent solution. He has 80 percent milk solution in stock, so he needs to dilute it with distilled water. How many ml of milk solution in stock and distilled water does chemist Rene needs?
Answer by stanbon(75887)

Rene, a chemist, needs to produce a 400 ml of 60 percent solution. He has 80 percent milk solution in stock, so he needs to dilute it with distilled water. How many ml of milk solution in stock and distilled water does chemist Rene needs?
------
Equation:
milk + milk = mild
0.80x + 0*(400-x) = 0.60*400
------
80x + 0 = 60*400
x = (3/4)(400)
x = 300 ml (amt of 80% milk needed)
400-x = 100 ml (amt of distilled water needed)
----------------
Cheers,
Stan H.
----------------